OpenStudy (anonymous):

How do I find the inverse(f^-1) of the function f(x)=1-2x-x^3? (Will give medals)

OpenStudy (anonymous):

The way I like to do it is by writing "f(x)" as "y"... y = (2x-3)/(x+1) And then switching your x's and your y's... x = (2y-3)/(y+1) and solving for y again. So... multiply both sides by (y+1): x(y+1) = (2y-3) xy + x = 2y - 3 get everything on one side (in this case, subtract (2y - 3) from both sides): xy + x - 2y + 3 = 0 Now, group your y's together, because remember, we want our answer to be y = ... xy - 2y + x + 3 = 0 Group and factor: y(x-2) + x + 3 = 0 Isolate y by moving all your x's to the one side (so subtract (X + 3) from both sides): y(x-2) = - x - 3 Divide both sides by (x - 2): y = - (x+3)/(x-2) And remember, the y in this equation is actually your inverse, so: The inverse is f '(x) = - (x+3)/(x-2) @Matthew071

OpenStudy (anonymous):

I understand the inverse concept I just need help finding the inverse of \[f(x)=1-2x-x^3\]